# V2X-M2C: Efficient Multi-Module Collaborative Perception with Two Connections

### v1

@@ Abstract
In this paper, we investigate improving the perception performance of autonomous vehicles through communication with other vehicles and road infrastructures. To this end, we introduce a collaborative perception model V2X-M2C, consisting of multiple modules, each generating inter-agent complementary information, spatial global context, and spatial local information. Inspired by the question of why most existing architectures are sequential, we analyze both the sequential and parallel connections of the modules. The sequential connection synergizes the modules, whereas the parallel connection independently improves each module. Extensive experiments demonstrate that V2X-M2C achieves state-of-the-art perception performance, increasing the detection accuracy by 8.00% to 10.87% and decreasing the FLOPs by 42.81% to 52.64%.
@@ 1 Introduction
One of the fundamental components of autonomous vehicles (AVs) is the ability to perceive various driving environments. With the advance of deep learning, perception systems of AVs have demonstrated effectiveness in various studies, including object detection [1, 14, 22] and segmentation[1, 23, 24]. However, the perception of a single vehicle alone still has limitations caused by occlusion and limited sensor range.
@@
To overcome these, multi-agent collaborative perception has been pivotal in enhancing perception across various environments. In particular, V2X (Vehicle-to-Everything) communications have enabled collaborative perception among heterogeneous agents such as vehicles and road infrastructure. Recognized methodologies such as CoBEVT [1], V2X-ViT [2], and Where2comm [3] have significantly improved the performance of 3D object detection.
The ongoing evolution of V2X collaborative perception necessitates continuous improvements. In particular, we develop an efficient model with manageable sizes and parameters. Consequently, our model can be implemented in lightweight hardware, becoming a more practical approach to collaborative perception. In addition, we propose a completely new concept for processing multi-agent features in collaborative perception. We have observed that most collaborative perception methods process multi-agent features sequentially. Because the sequential architecture handles the features as a whole and does not consider the individual features, it cannot always achieve the desired performance. By introducing and analyzing a parallel architecture, we open up new promising architecture directions in collaborative perception.
@@
Figure 1: About this paper. (a) and (b) present two connections, and the attention weight resulted in M2C-S and M2C-P (details in Section 4.3). (c) V2X-M2C holds state-of-the-art performance with a reduced number of model parameters (details in Section 4.2)
In this paper, we introduce V2X-M2C (V2X-Multi Modules with Two Connections), a framework designed to elevate effectiveness and efficiency based on parallel and sequential connections. Figure 2 illustrates two connections within holistic architecture, and Figure 3 depicts three key sub-modules: i) agent-wise attention module (A-Att) with heterogeneous relative pose encoding (HRPE) that interprets interactions between multiple agents; ii) spatial-wise attention module (S-Att) that captures per-agent spatial global context; iii) heterogeneity-reflected convolution module (H-Conv) that generates spatial local information depending on agent types. We summarize our contributions as follows.
  • We propose the collaborative perception model, V2X-M2C for 3D object detection. V2X-M2C is divided into two models: M2C-S connects the sub-modules in series, and M2C-P connects them in parallel. We consider multiple heterogeneous agents, making our model expandable to V2X scenarios. Our two models achieve state-of-the-art effectiveness, efficiency, and robustness to noise.
  • We perform the qualitative comparison between the two connection types regarding how each sub-module influences the others. (Section 4.3)
@@
  • We utilize a channel compression layer (CCL) to make the overall fusion module lightweight, thereby significantly enhancing model efficiency while maintaining performance.
  • We design a heterogeneity-reflected convolution module (H-Conv) to consider multiple heterogeneous agents with a low number of parameters and low computation complexity.
@@ 2 Related Work
Collaborative Perception. Intermediate fusion is the main technique that enables collaborative perception, which balances performance and communication demands. Previous studies obtain and share information using knowledge distillation [6] , multi-module [1, 2, 5], and spatial confidence-aware communication [3] . They have proven effective in modeling heterogeneous data. However, their potential to reduce computational complexity and capture local context remains understudied. In this paper, we design the two different connection models using multi-modules to ensure model efficiency while outperforming the previous methods.
Computer vision. The introduction of the convolution layer has revolutionized computer vision; for example, AlexNet [7], VggNet [8] , ResNet [9], etc. The convolution layer effectively captures local features while requiring fewer parameters than a fully connected layer. However, the convolution layer has difficulty capturing the relationships between distant features. The Vision Transformer(ViT) [10] leverages self-attention mechanisms to capture such global relationships effectively, yet at high computation complexity and with the challenge of maintaining translation equivariance. Several approaches, such as Swin Transformer [11] , Pooling-based Vision Transformer [12] , Neighborhood Attention Transformer [13], have been introduced to solve these challenges. Similar to these approaches, our approach takes full advantage of convolution and transformer to capture local information and global context, respectively.
Figure 2: Overview of our proposed collaborative perception system. Our model consists of five steps: metadata sharing, feature extraction, feature sharing, fusion module, and detection head. The details of each component are discussed in Section 3. (a) shows preprocessing, (b) and (c) depict M2C-S and M2C-P fusion of multi-modules.
@@ 3 Method
The overall architecture of V2X-M2C is shown in Figure 2. Depending on the connection type, the input of the sub-modules and the overall process differ. To distinguish it, we use $\mathrm{Conn.}\in\{\mathbf{S},\mathbf{P}\}$, where $\mathbf{S}$ indicates the M2C-S and $\mathbf{P}$ the M2C-P. In this section, we introduce the five main components: 1) metadata sharing, 2) feature extraction, 3) feature sharing, 4) fusion module, and 5) detection head.
@@ 3.2 Feature Extraction
Our proposed model is based on intermediate fusion, which shares features extracted from raw point-cloud data. In V2X-M2C, the feature extraction mainly consists of 1)PointPillars and 2)Sparse Resnet Backbone.
PointPillars. PointPillars[14] splits point clouds into vertical columns. This enables PointPillars to use less memory and become faster than other voxel-based approaches, such as [15]. Also, the results of PointPillars are 2D pseudo-images. Thus, it is proper to apply a 2D convolution layer. To achieve efficient extraction, we employ PointPillars to convert point clouds of all agents to 2D pseudo-images.
Sparse Resnet Backbone. Sparsity is an important property of point clouds. The works [16, 25] suggest sub-manifold sparse convolution and sparse convolution, both of which are more effective and efficient for extracting features from sparse data. They extract feature $\mathbf{F}_{l}\in\mathbb{R}^{H\times W\times C}$ from the $l$-th agent’s 2D pseudo-image where $H$ is the height, $W$ the width, and $C$ the channels. All agents use the same backbone parameters. We deploy the Sparse Resnet(SpRes) backbone [28, 29, 30], which reduces memory usage and works more effectively than conventional dense convolution.
@@ 3.3 Feature Sharing
$\mathbf{F}^{\prime}_{l}=f_{\mathrm{encoder}}(\mathbf{F}_{l}),\quad l=1,...,L$
$\mathbf{\bar{F}}_{l}=f_{\mathrm{decoder}}(\mathbf{F}^{\prime}_{l}),\quad l=1,...,L.$
Transformation. The ego agent transforms the decoded features into the ego agent coordinate. Also, features outside the detection range are cropped to delete unnecessary information. After this process, we use a spatial-temporal correction module (STCM) [2] to compensate for the time delay caused by communication latency. We denote by $\Gamma(\cdot)$ the sequential process of transform, crop, and STCM. By concatenating $\Gamma(\mathbf{\bar{F}_{l}})$ along the agent dimension, the multi-agent feature $\mathbf{M^{0}}\in\mathbb{R}^{L\times H\times W\times C}$ is generated. That is,
$\mathbf{M^{0}}=\underset{l\in[1,L]}{\parallel}\Gamma(\mathbf{\bar{F}}_{l})$
@@ 3.4 Fusion Module
Figure 3: Sub-modules in V2X-M2C. (a) Agent-wise Attention (A-Att) sub-module with HRPE. (b) Spatial-wise Attention (S-Att) sub-module. (c) Heterogeneity-reflected Convolution (H-Conv) sub-module discussed in Sec. 3.4.
Our fusion module in V2X-M2C mainly consists of three sub-modules: agent-wise attention (A-Att), spatial-wise attention (S-Att), and heterogeneity-reflected convolution (H-Conv). Each sub-module generates inter-agent complementary information, spatial global context, and spatially detailed information, respectively. We refer to these as the intent of each sub-module. To prevent the model from being too heavy due to two transformer sub-modules (A-Att, S-Att), we use a channel-wise compression layer (CCL) to reduce the input feature size. This is one of the key components that makes our model efficient while maintaining state-of-the-art performance. Also, we propose two different module connections, sequential and parallel.
Channel Compression Layer (CCL). To reduce the input feature size in a simple and information-maintained manner, we design CCL as one fully connected (FC) layer that compresses the feature channel-wise. The input to the fusion module is the multi-agent feature $\mathbf{M}^{0}$, and the input at the $d$-th depth is $\mathbf{M}^{d}\in\mathbb{R}^{L\times H\times W\times C},d=1,...,D-1$ where $D$ is the maximum depth of the fusion module. Then, the output $\mathbf{M}^{\prime}$ of CCL at the $d$-th depth for each $d=0,...,D-1$ is given by
@@
CCL plays two roles in M2C-P. First, M2C-P generates the output by concatenating four intents along the channel. CCL compresses the feature channel size from $C$ to $C/4$ to match the input and output size. Second, from the second depth of the fusion module, CCL fuses the input feature channel-wise, and CCL actively improves the deficiency of each intent by adopting the strength of the others.
@@
According to the $l$-th agent type $t_{l}\in\{\mathrm{I},\mathrm{V}\}$, we define the period $w_{j}$ as the inverse of the constant hyperparameter $\tau\in\mathbb{R}$ to some powers of integer $j\in[0,C/16-1]$. In particular,
$\omega_{j}=\begin{cases}1/(\tau^{2j+1}),&t_{l}=\mathrm{I}\\ 1/(\tau^{2j}),&t_{l}=\mathrm{V}.\\ \end{cases}$
Using the approach in [26], we make fixed initialized encoding values $\mathbf{p}\in\mathbb{R}^{C/4}$ of the relative angle $\theta$ and distance $d$ between the ego and aux agents as
@@
where * is the multiplication.
After the initialization, the fixed encoding value passes the linear projection $f_{\mathrm{I}}$ or $f_{\mathrm{V}}$ depending on the agent type. It is added to the $l$-th compressed input feature $\mathbf{M}^{\prime}_{l}\in\mathbb{R}^{H\times W\times C/4}$, and then, the encoded feature $\mathbf{M}^{\mathrm{pos}}$ concatenates the added feature along the agent dimension. That is,
$\mathbf{M}^{\mathrm{pos}}=\underset{l\in[1,L]}{\parallel}(\mathbf{M}^{\prime}_{l}+f_{t_{l}}(\mathbf{p})).$
Agent-wise Attention Module (A-Att). We use the vanilla attention module available in the code of [2], which we denote by $\mathrm{CAV\_ATT}^{d}$ at the $d$-th depth. This module does not distinguish agent type and utilizes a multi-head self-attention transformer. For each connection $\mathrm{Conn.}$, the output ${}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathbf{AA}}$ of A-Att at the $d$-th depth is given by
${}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathbf{AA}}=\begin{cases}\mathrm{CAV\_ATT}^{1}(\mathbf{M}^{\mathrm{pos}})+\mathbf{M}^{\prime},&\quad d=1\\ \mathrm{CAV\_ATT}^{d}(\mathbf{M}^{\prime})+\mathbf{M}^{\prime},&\quad d\neq 1.\end{cases}$
@@
Spatial-wise Attention Module (S-Att). We use Dilated Neighborhood Attention Transformer (DiNAT)[17] to interpret a global context. DiNAT maintains translation equivariance and reduces computation burden by resembling the convolution operation. Also, it efficiently widens the receptive fields by using the dilation rate. We use one block of the original DiNAT module and make it lighter to enhance the overall efficiency of the model. The output ${}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathbf{SA}}$ of S-Att is determined by
${}^{{\mathbf{S}}}\mathbf{M}^{\mathbf{SA}}$ $=\mathrm{DiNAT}({}^{{\mathbf{S}}}\mathbf{M}^{\mathbf{AA}})+{}^{{\mathbf{S}}}\mathbf{M}^{\mathbf{AA}},$
${}^{{\mathbf{P}}}\mathbf{M}^{\mathbf{SA}}$ $=\mathrm{DiNAT}(\mathbf{M}^{\prime})+\mathbf{M}^{\prime}.$
Heterogeneity-reflected Convolution Module (H-Conv). To consider multiple heterogeneous agents, most prior works use transformer-based approaches [2, 32, 33]. On the other hand, we use convolution layers to handle multiple heterogeneous agents and prevent our model from being computationally heavy. H-Conv uses two pipelines, which have the same architecture but are distinguished by agent types. Each pipeline has two convolution layers with residual connections. In each pipeline, the $l$-th agent input feature to H-Conv is first divided into ${}^{\mathrm{Conn.}}\mathbf{M}^{\mathrm{I}}_{l}$ and ${}^{\mathrm{Conn.}}\mathbf{M}^{\mathrm{V}}_{l}$ depending on the agent types. Then, these are concatenated along the agent dimension, generating ${}^{\mathrm{Conn.}}\mathbf{M}^{\mathrm{I}}$ (${}^{\mathrm{Conn.}}\mathbf{M}^{\mathrm{V}}$). Then, ${}^{\mathrm{Conn.}}\mathbf{M}^{\mathrm{I}}$ (${}^{\mathrm{Conn.}}\mathbf{M}^{\mathrm{V}}$) is passed through the convolution operation $f_{\mathrm{conv}}^{\mathrm{I}}(\cdot)$ ($f_{\mathrm{conv}}^{\mathrm{V}}(\cdot)$). The output ${}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathbf{HC}}$ of H-Conv is determined by
${}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathbf{HC}}=f_{\mathrm{conv}}^{\mathrm{I}}({}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathrm{I}})+f_{\mathrm{conv}}^{\mathrm{V}}({}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathrm{V}})$
where
${}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathrm{I}}=\underset{l\in[1,L]}{\parallel}{}^{{\mathrm{Conn.}}}\mathbf{M}_{l}^{\mathrm{I}},$ ${}^{{\mathrm{Conn.}}}\mathbf{M}^{\mathrm{V}}=\underset{l\in[1,L]}{\parallel}{}^{{\mathrm{Conn.}}}\mathbf{M}_{l}^{\mathrm{V}}$
${}^{{\mathbf{S}}}\mathbf{M}^{t}_{l}=\begin{cases}{}^{{\mathbf{S}}}\mathbf{M}^{\mathbf{SA}}_{l},&\quad t=t_{l}\\ 0,&\quad t\neq t_{l}\end{cases}\ ,\quad$ ${}^{{\mathbf{P}}}\mathbf{M}^{t}_{l}=\begin{cases}\mathbf{M}^{\prime}_{l},&\quad t=t_{l}\\ 0,&\quad t\neq t_{l}.\end{cases}$
@@
Sequential/Parallel Connection (M2C-S/M2C-P). M2C-S puts the three sub-modules (A-Att, S-Att, and H-Conv) in series and, at the end, uses the channel expand layer (CEL), $f_{\mathrm{CEL}}:\mathbb{R}^{C/4}\to\mathbb{R}^{C}$, to restore the original channel size of $C$. M2C-P concatenates the intents of the three sub-modules and the compressed original feature $M^{\prime}$. Then, the output of $d$-th depth fusion module $\mathbf{M}^{d}$ is generated by applying a multilayer perceptron (MLP) with the residual connection. That is,
$\mathbf{M}^{d}=\mathrm{MLP}(\mathbf{M}^{\mathbf{O}})+\mathbf{M}^{\mathbf{O}},\quad d=1,...,D$
where
$\mathbf{M}^{\mathbf{O}}=\begin{cases}f_{\mathrm{CEL}}\left({}^{{\mathbf{S}}}\mathbf{M}^{\mathbf{HC}}\right),\quad&\mathrm{Conn.}=\mathbf{S}\\ {}^{{\mathbf{P}}}\mathbf{M}^{\mathbf{AA}}\parallel{}^{{\mathbf{P}}}\mathbf{M}^{\mathbf{SA}}\parallel{}^{{\mathbf{P}}}\mathbf{M}^{\mathbf{HC}}\parallel\mathbf{M}^{\prime},\quad&\mathrm{Conn.}=\mathbf{P}.\end{cases}$
Depending on the connection type, the learning tendency of the sub-modules is different. The property of M2C-S is that the input of one sub-module is the output of the other. This dependency makes the sub-modules learn to enhance their performance as a whole. M2C-P concatenates all the intents and applies MLP. Then, the output $\mathbf{M}^{d}$ is applied to the $(d+1)$-th depth fusion module as the input feature. The FC layers in MLP and CCL learn to emphasize the strength of each intent to compensate for the deficiency of the others. In other words, the sub-modules learn to strengthen what they do best to generate rich information for detection. Our detailed analysis of these learning tendencies is presented in Section 4.3.

### v2

# ParCon: Noise-Robust Collaborative Perception via Multi-module Parallel Connection

@@ Abstract
In this paper, we investigate improving the perception performance of autonomous vehicles through communication with other vehicles and road infrastructures. To this end, we introduce a novel collaborative perception architecture, called ParCon, which connects multiple modules in parallel, as opposed to the sequential connections used in most other collaborative perception methods. Through extensive experiments, we demonstrate that ParCon inherits the advantages of parallel connection. Specifically, ParCon is robust to noise, as the parallel architecture allows each module to manage noise independently and complement the limitations of other modules. As a result, ParCon achieves state-of-the-art accuracy, particularly in noisy environments, such as real-world datasets, increasing detection accuracy by 6.91%. Additionally, ParCon is computationally efficient, reducing floating-point operations (FLOPs) by 11.46%.
@@ 1 Introduction
One of the fundamental components of autonomous vehicles (AVs) is the ability to perceive various driving environments. With the advance of deep learning, perception systems of AVs have demonstrated effectiveness in various studies, including object detection (Liu et al. [2023]; Hu et al. [2023]; Kumar et al. [2024]) and segmentation (Zhang et al. [2023]; Xu et al. [2023]; Wu et al. [2024]). However, the perception of a single vehicle alone still has limitations caused by occlusion and limited sensor range.
@@
To overcome these, multi-agent collaborative perception has been pivotal in enhancing perception across various environments. In particular, V2X (Vehicle-to-Everything) communications have enabled collaborative perception among heterogeneous agents such as vehicles and road infrastructure. Recognized methodologies, such as CoBEVT (Xu et al. [2022a]), V2X-ViT (Xu et al. [2022b]), and Where2comm (Hu et al. [2022]), have significantly improved the performance of 3D object detection.
One of the major challenges in collaborative perception is that the ego agent must communicate with surrounding agents, inevitably introducing noise during communication. However, most V2X collaborative perception models employ a sequential architecture to connect agent-wise and spatial-wise modules (see Figure 1 (a)), which tends to amplify the impact of noise as it propagates through the entire process. To solve this problem, parallel architectures have been adopted in various fields (Kim et al. [2018]; Kang et al. [2023]). When applied to V2X systems, parallel architectures offer significant benefits, including improved robustness to communication noise and enhanced performance. To leverage these advantages, we propose ParCon, a novel V2X collaborative perception model for 3D object detection that features parallel connections as shown in Figure 1 (b).
@@
Figure 1: About this paper. (a) and (b) present two architectures. (c) and (d) ParCon holds state-of-the-art performance with a reduced number of model parameters and under various noises (details in Section 4.2)
In this paper, we introduce our novel model, ParCon, and present extensive simulation results on various collaborative perception datasets. We demonstrate that ParCon outperforms other state-of-the-art (SOTA) approaches in detection accuracy and exhibits substantially increased robustness to communication noise (see Figure 1 (c), (d)). In particular, this robustness enables the model to achieve significantly improved detection accuracy in noisy environments, including real-world datasets. Furthermore, ParCon uses fewer parameters and is more computationally efficient than other SOTA approaches (see Figure 1 (c)).
We summarize our contributions as follows.
  • We propose a novel parallel connection-based collaborative perception model, called ParCon, for 3D object detection. Our model achieves state-of-the-art accuracy, computation efficiency, and robustness to noise.
  • We demonstrate the advantages of parallel connection. Namely, we highlight the robustness of our model to communication noise by analyzing it at different levels of communication latency, heading errors, and localization errors.
@@
  • We utilize a channel compression layer (CCL) to fuse various types of information, including agent-wise and spatial-wise, global and local information, and make the overall fusion module lightweight, thereby significantly enhancing model efficiency.
@@ 2 Related Work
Collaborative Perception. Intermediate fusion, which fuses compressed features rather than raw sensor data, is the primary technique enabling collaborative perception to balance performance with communication demands. Previous studies obtain and share information using knowledge distillation (Mehr et al. [2019]), multi-module (Xu et al. [2022a]; Xu et al. [2022b]; Yang et al. [2023]), and spatial confidence-aware communication (Hu et al. [2022]). Also, to reduce the impact of noise, many studies have attempted to develop robust models, such as V2Vnet(Wang et al. [2020]), V2X-ViT(Xu et al. [2022b]), FeaCo(Gu et al. [2023]), and CoAlign(Lu et al. [2023]). To the best of our knowledge, these models fuse the information in series, and the potential of using a parallel connection has not been explored. In this paper, we design a new collaborative perception model that connects the information in parallel to enhance model efficiency and noise robustness.
Computer vision. The introduction of convolutional layers has revolutionized the field of computer vision; for example, AlexNet (Krizhevsky et al. [2012]), VggNet (Simonyan & Zisserman [2014]), ResNet (He et al. [2016]), etc. Convolutional layers effectively capture local features while requiring fewer parameters than fully connected layers. However, they struggle to capture relationships between distant features. Vision Transformer (ViT) (Dosovitskiy et al. [2020]) leverages self-attention mechanisms to effectively capture global relationships, but at the cost of high computational complexity and challenges in maintaining translation equivariance. Several approaches, such as Swin Transformer (Liu et al. [2021]) and Neighborhood Attention Transformer (Hassani et al. [2023]), have been introduced to solve these challenges. Similar to these approaches, our approach fully exploits the strengths of both convolution and transformer architectures to capture local features and global context, respectively.
Figure 2: Overview of our proposed collaborative perception system. Our model consists of five steps: metadata sharing, feature extraction, feature sharing, fusion module, and detection head. The details of each component are discussed in Section 3.
@@ 3 Method
The overall architecture of ParCon is shown in Figure 2. In this section, we introduce the five main components: 1) metadata sharing, 2) feature extraction, 3) feature sharing, 4) fusion module, and 5) detection head.
@@ 3.2 Feature Extraction
Most V2X collaborative perception models are based on intermediate fusion, which shares features extracted from raw point-cloud data. In ParCon, the feature extraction mainly consists of 1) PointPillars and 2) Sparse Resnet Backbone.
PointPillars. PointPillars (Lang et al. [2019]) splits point clouds into vertical columns. This enables PointPillars to use less memory and become faster than other voxel-based approaches, such as (Zhou & Tuzel [2018]). Also, the results of PointPillars are 2D pseudo-images. Thus, it is proper to apply a 2D convolution layer. To achieve efficient extraction, we employ PointPillars to convert point clouds of all agents to 2D pseudo-images.
Sparse Resnet Backbone. Sparsity is an important property of point clouds. The works (Graham et al. [2018]; Yan et al. [2018]) suggest sub-manifold sparse convolution and sparse convolution, both of which are more effective and efficient for extracting features from sparse data. They extract feature $\mathbf{F}_{l}\in\mathbb{R}^{H\times W\times C}$ from the $l$-th agent’s 2D pseudo-image where $H$ is the height, $W$ the width, and $C$ the channels. All agents use the same backbone parameters. We deploy the Sparse Resnet (SpRes) backbone (Shi et al. [2022]; Yin et al. [2021]; Zhu et al. [2019]), which reduces memory usage and works more effectively than conventional dense convolution.
@@ 3.3 Feature Sharing
$\mathbf{F}^{\prime}_{l}=f_{\mathrm{encoder}}(\mathbf{F}_{l}),\quad l=1,...,L,$ $\mathbf{\bar{F}}_{l}=f_{\mathrm{decoder}}(\mathbf{F}^{\prime}_{l}),\quad l=1,...,L.$
Transformation. The ego agent transforms the decoded features into the ego agent coordinate. Also, features outside the detection range are cropped to delete unnecessary information. After this process, we use a spatial-temporal correction module (STCM) used in Xu et al. [2022b] to compensate for the time delay caused by communication latency. We denote by $\Gamma(\cdot)$ the sequential process of transform, crop, and STCM. By concatenating $\Gamma(\mathbf{\bar{F}_{l}})$ along the agent dimension, the multi-agent feature $\mathbf{M^{0}}\in\mathbb{R}^{L\times H\times W\times C}$ is generated. That is,
$\mathbf{M^{0}}=\underset{l\in[1,L]}{\parallel}\Gamma(\mathbf{\bar{F}}_{l}),$
@@ 3.4 Fusion Module
Figure 3: Sub-modules in ParCon. (a) Agent-wise Attention (A-Att) sub-module with HRPE. (b) Spatial-wise Attention (S-Att) sub-module. (c) Spatial-wise Convolution (S-Conv) sub-module. These modules are discussed in Section 3.4.
The overall architecture of the fusion module is described in Figure 3 (b). Our fusion module in ParCon mainly consists of three sub-modules: agent-wise attention (A-Att), spatial-wise attention (S-Att), and spatial-wise convolution layer (S-Conv). Each sub-module generates inter-agent complementary information, spatial global context, and spatially detailed information, respectively. We refer to the output of each sub-module as the intent of each sub-module. To prevent the model from being too heavy due to two transformer sub-modules (A-Att, S-Att), we use a channel-wise compression layer (CCL) to generate a compressed feature which is compressed the input feature channel-wise. This is one of the key components that makes our model efficient while maintaining state-of-the-art performance. After intents are generated from sub-modules, the output feature concatenates all intents and the compressed feature.
Channel Compression Layer (CCL). To reduce the input feature size in a simple and information-maintained manner, we design CCL as one fully connected (FC) layer that compresses the feature channel-wise. The input to the fusion module is the multi-agent feature $\mathbf{M}^{0}$, and the input at the $d$-th depth is $\mathbf{M}^{d}\in\mathbb{R}^{L\times H\times W\times C},d=1,...,D-1$ where $D$ is the maximum depth of the fusion module. Then, the compressed feature $\mathbf{M}^{\prime}$, output of CCL, at the $d$-th depth for each $d=0,...,D-1$ is given by
@@
Although CCL is a simple FC layer, it plays two important roles in ParCon. First, by compressing the feature channel size from $C$ to $C/4$, CCL significantly reduces the number of parameters and GFLOPs in ParCon. Second, because CCL mixes various intents to generate $\mathbf{M^{\prime}}$, it contributes to improved performance by compensating for the deficiencies of individual modules and enhancing overall robustness.
@@
According to the $l$-th agent type $t_{l}\in\{\mathrm{I},\mathrm{V}\}$, we define the period $w_{j}$ as the inverse of the constant hyperparameter $\tau\in\mathbb{R}$ to some powers of integer $j\in[0,C/16-1]$. In particular, $\omega_{j}=1/(\tau^{2j+1})$ if the agent type of $l$ is infrastructure, and $\omega_{j}=1/(\tau^{2j})$ if the agent type of $l$ is vehicle. Using the approach in Piergiovanni et al. [2023], we make fixed encoding values $\mathbf{p}\in\mathbb{R}^{C/4}$ of the relative angle $\theta$ and distance $d$ between the ego and aux agents as
@@
where * is the multiplication. The fixed encoding value $\mathbf{p}$ is added to the $l$-th compressed input feature $\mathbf{M}^{\prime}_{l}\in\mathbb{R}^{H\times W\times C/4}$, and then, the encoded feature $\mathbf{M}^{\mathrm{pos}}\in\mathbb{R}^{L\times H\times W\times C/4}$ concatenates the added feature along the agent dimension. That is, $\mathbf{M}^{\mathrm{pos}}=\underset{l\in[1,L]}{\parallel}(\mathbf{M}^{\prime}_{l}+\mathbf{p}).$
Agent-wise Attention Module (A-Att). As shown in Figure 3 (a), we use the vanilla attention module available in the code of Xu et al. [2022b], which we denote by $\mathrm{CAV\_ATT}^{d}$ at the $d$-th depth. This module does not distinguish agent type and utilizes a multi-head self-attention transformer. The output ${}\mathbf{M}^{\mathbf{AA}}$ of A-Att at the $d$-th depth is given by
$\mathbf{M}^{\mathbf{AA}}=\begin{cases}\mathrm{CAV\_ATT}^{1}(\mathbf{M}^{\mathrm{pos}})+\mathbf{M}^{\prime},&\quad d=1\\ \mathrm{CAV\_ATT}^{d}(\mathbf{M}^{\prime})+\mathbf{M}^{\prime},&\quad d\neq 1.\end{cases}$
@@
Spatial-wise Attention Module (S-Att). We use Dilated Neighborhood Attention Transformer (DiNAT) (Hassani & Shi [2022]) to interpret a global context. DiNAT maintains translation equivariance and reduces computation burden by resembling the convolution operation. Also, it efficiently widens the receptive fields by using the dilation rate. As shown in Figure 3 (b), we use one block of the original DiNAT module and make it lighter by reducing the depth, kernel size, and dilation value of the original DiNAT to enhance the overall efficiency of the model. The output $\mathbf{M}^{\mathbf{SA}}$ of S-Att is determined by
$\mathbf{M}^{\mathbf{SA}}$ $=\mathrm{DiNAT}(\mathbf{M}^{\prime})+\mathbf{M}^{\prime}.$
Spatial-wise Convolution Module (S-Conv). We use convolution layers to capture the spatial local information. As shown in Figure 3 (c), S-Conv uses three blocks of convolution layers with batch normalization and activate function ReLU. Also, we apply residual connections only to the first and second blocks while omitting them in the third block. The output $\mathbf{M}^{\mathbf{SC}}$ of S-Conv is determined by
$\mathbf{M}^{\mathbf{SC}}=\mathrm{S\_Conv}(\mathbf{M}^{\prime})$
@@
Parallel Connection. ParCon concatenates the intents of the three sub-modules and the compressed original feature $\mathbf{M^{\prime}}$. Then, the output of $d$-th depth fusion module $\mathbf{M}^{d}$ is generated by applying a multilayer perceptron (MLP) with the residual connection. That is,
$\mathbf{M}^{\mathbf{O}}=\mathbf{M}^{\mathbf{AA}}\parallel\mathbf{M}^{\mathbf{SA}}\parallel\mathbf{M}^{\mathbf{SC}}\parallel\mathbf{M}^{\prime},$ $\mathbf{M}^{d}=\mathrm{MLP}(\mathbf{M}^{\mathbf{O}})+\mathbf{M}^{\mathbf{O}},\quad d=1,...,D.$